OpenStudy (anonymous):

I must use the formula A=Pe^rt. I have been given P=5,000 and the amount after 10 years ($5665.74). I have to find the annual % rate and the time it takes to double. Please leave detialed step by step instructions (all in one comment.)

OpenStudy (anonymous):

A=Pe^rt P=5000 A=5665.74 e=2.718281 t=10 Find r.

OpenStudy (anonymous):

Its a pretty simple question and I think you can solve it from here :3

OpenStudy (anonymous):

So I have everything all plugged into where It should go, but I was gone for the lesson when my teacher taught us how to do it, so I have no idea where to start after that. (do I divide by 5,000 on each side? If so what do I do next?)

OpenStudy (hlares):

Yep! If you have the equation set up like 5,665.74=5,000*e^(10r), then you will begin by dividing both sides by 5,000. After that, convert the x=e^c equation that you have into lnx=c, then solve for r.